OpenStudy (anonymous):

find the roots of the quadratic equation x^2 -8x=9 by completing the square

OpenStudy (lynfran):

\[x ^{2}-8x=9\]\[x ^{2}-8x-9=0\]....can you tell me what 2numbers when multiple =-9 and when added together=-8

OpenStudy (lynfran):

@ginagasko ??

OpenStudy (lynfran):

ok -9*1=-9 and -9+1=-8 so....we have and equation like this\[x ^{2}-9x+x-9=0\]we then factor out the like terms we get\[x(x-9)+1(x-9)=0\]so we now how \[(x+1)(x-9)=0\]now im gonna let the part for you to complete...hint set each bracket to zero and solve for x